zarkam21 (zarkam21):

Write the equations for two different lines that intersect at (-2,3)

OpenStudy (anonymous):

y1 = -2x + 7 y2 = x + 5 y1 = -2(-2) + 7 = 3 y2 = -2 + 5 = 3

zarkam21 (zarkam21):

Which two is it?

OpenStudy (anonymous):

I belive it is the top two

OpenStudy (alexandervonhumboldt2):

these are the same equation. in next paragtah he showed that they result in same nimber.

OpenStudy (alexandervonhumboldt2):

y1 = -2x + 7 y2 = x + 5 is the pair of equation

OpenStudy (alexandervonhumboldt2):

here he showes that they result both in 3 y1 = -2(-2) + 7 = 3 y2 = -2 + 5 = 3

zarkam21 (zarkam21):

OKay thanks could you help me with the slope of both

OpenStudy (alexandervonhumboldt2):

y1 = -2x + 7 slope is -2 y2 = x + 5 slope is 1 y=ax+b where a is slope.

zarkam21 (zarkam21):

Thank you soo much